# Competition of Mechanisms: Tracing How Language Models Handle Facts and Counterfactuals

### v1

@@ Appendix B Experimental Details
Figure S1: Layer-wise Position Analysis of Relevant Tokens in GPT-2-small. The figure presents the logit values for two pertinent tokens across various positions and layers. The left panel illustrates the logit values for the factual token $t_{\mathrm{fact}}$, while the right panel illustrates the logit values for the counterfactual token $t_{\mathrm{cofa}}$.
Figure S2: Attribution of Logit Differences to Attention and MLP Blocks. This figure depicts the role of Attention and MLP Blocks in influencing the logit difference $\Delta_{\mathrm{cofa}}=\text{Logit}(t_{\mathrm{cofa}})-\text{Logit}(t_{\mathrm{fact}})$ at $W_{U}\mathbf{a_{N}}^{l}$ for the final position of the input across various layers. It is observed that the Attention Blocks have a more pronounced impact compared to the MLP Blocks. Additionally, the right panel indicates a contribution from the MLP Blocks at the last layers towards the factual token $t_{\mathrm{fact}}$.
Figure S3: Attention Pattern for Relevant Attention Heads. The panel illustrates the attention patterns of relevant heads for the last position, demonstrating consistent attention to the attribute position by both red and blue heads.
@@
Figure S4: Direct Contribution of Attention Heads. The figure displays the direct contribution of all heads in Pythia-6.9b to the logit difference $\Delta_{\text{cofa}}$ with heads favoring $t_{\mathrm{fact}}$ highlighted in red and those favoring $t_{\mathrm{cofa}}$ in blue.
Appendix D Other Experiment for GPT-2
@@
Figure S5: Rank of Target Tokens for Attribute Position Across Layers in GPT-2. This figure depicts the trend where the logit rank for the factual token $t_{\mathrm{fact}}$ decreases while the rank for the counterfactual token $t_{\mathrm{cofa}}$ increases at the attribute position. In the concluding layers, this pattern is evident as $t_{\mathrm{fact}}$ typically secures a lower rank, in contrast to $t_{\mathrm{cofa}}$, which shows an upward trajectory in rank. However, it is important to note that $t_{\mathrm{cofa}}$’s rank consistently remains lower than that of $t_{\mathrm{fact}}$.
@@
Figure S6: Attention Pattern of Significant Heads. This figure illustrates the comprehensive attention pattern of heads substantially influencing $\Delta_{t_{\mathrm{cofa}}}$. Notably, a similar pattern emerges for both heads favoring $t_{\mathrm{cofa}}$ (depicted in blue) and those favoring $t_{\mathrm{fact}}$ (illustrated in red), particularly in the attention edge between the attribute and the final position.

### v2

@@ A.1 Macroscopic Inspection across Layers and Token Positions
Figure 7: Layer-wise Position Analysis of Relevant Tokens in GPT-2-small. The figure presents the logit values for two pertinent tokens across various positions and layers. The left panel illustrates the logit values for the factual token $t_{\mathrm{fact}}$, while the right panel illustrates the logit values for the counterfactual token $t_{\mathrm{cofa}}$.
A.2 Intermediate Inspection of Attention and MLP Blocks
This subsection exposes the contributions of Attention and MLP Blocks to the differences in logit values across layers within Pythia-6.9b. Figure 8 explores how these components influence the computation of logits for two tokens, represented as the difference $\Delta_{\mathrm{cofa}}=\text{Logit}(t_{\mathrm{cofa}})-\text{Logit}(t_{\mathrm{fact}})$ at the final position of the input. The analysis specifically highlights the distinct effects of these blocks at different stages of the model’s operation.
Figure 8: Attribution of Logit Differences to Attention and MLP Blocks. delineates the differential impact of Attention and MLP Blocks on logit values at the terminal input position. The attention mechanism is shown to predominantly influence early layer processing in the left panel, while the right panel details the increased contribution of MLP Blocks to the factual token’s logits in the concluding layers, illustrating the dynamic interplay between these fundamental neural network elements.
A.3 Microscopic Inspection of Individual Attention Heads
Figure 9 quantifies the direct contributions of all attention heads to the difference in logit values, labeled as $\Delta_{\text{cofa}}$. It specifically identifies heads that preferentially enhance the logits for $t_{\mathrm{fact}}$ (shown in blue) versus those favoring $t_{\mathrm{cofa}}$ (depicted in red), offering insights into how attention mechanisms differentially prioritize token attributes.
@@ A.3 Microscopic Inspection of Individual Attention Heads
Figure 9: Direct Contribution of Attention Heads. The figure displays the direct contribution of all heads in Pythia-6.9b to the logit difference $\Delta_{\text{cofa}}$ with heads favoring $t_{\mathrm{fact}}$ highlighted in blue and those favoring $t_{\mathrm{cofa}}$ in red.
Figure 10 presents the attention patterns of the relevant attention heads at the last token position. It shows the consistent pattern of the relevant heads, with a consistent focus on the attribute position.
Figure 10: Attention Pattern for Relevant Attention Heads. The panel illustrates the attention patterns of relevant heads for the last position, demonstrating consistent attention to the attribute position by both red and blue heads.
Appendix B Other Experiment for GPT-2
B.1 Ranks Analysis in the Last Position
We provide additional information in Figure 11 mapping the logits to ranks of the tokens, and find that the rank of $t_{\mathrm{cofa}}$ in the projected logit distribution remains very low: $t_{\mathrm{cofa}}$ is among the 20 most likely tokens in the first five layers and between the 20th and the 70th in the last part of the network.
@@ B.1 Ranks Analysis in the Last Position
Figure 11: Rank of Target Tokens for Attribute Position Across Layers in GPT-2. This figure depicts the trend where the logit rank for the factual token $t_{\mathrm{fact}}$ decreases while the rank for the counterfactual token $t_{\mathrm{cofa}}$ increases at the attribute position. In the concluding layers, this pattern is evident as $t_{\mathrm{fact}}$ typically secures a lower rank, in contrast to $t_{\mathrm{cofa}}$, which shows an upward trajectory in rank. However, it is important to note that $t_{\mathrm{cofa}}$’s rank consistently remains lower than that of $t_{\mathrm{fact}}$.
B.2 Attention Pattern of Relevant Attention Heads
Figure 12 shows the full attention pattern for the relevant attention heads, as identified in Section 6. It is show as the attention pattern is similar between all the relevant attention heads, independently if the heads is favoring $t_{\mathrm{fact}}$ or $t_{\mathrm{cofa}}$.
@@ B.2 Attention Pattern of Relevant Attention Heads
Figure 12: Attention Pattern of Significant Heads. This figure illustrates the comprehensive attention pattern of heads substantially influencing $\Delta_{t_{\mathrm{cofa}}}$. Notably, a similar pattern emerges for both heads favoring $t_{\mathrm{cofa}}$ (depicted in red) and those favoring $t_{\mathrm{fact}}$ (illustrated in blue), particularly in the attention edge between the attribute and the final position.